hari borrowed Rs 12600 from a moneylender at 15% per annum simple interest. After 3 years, he paid Rs 7070 and gave a goat to clear off the debt. what is the cost of the goat.
Answer by mananth(16948)

SI = P * R * N/100
=12600*3*15/100
=126*3*15
=5670
Interest = 5670 + Principal 12600=17670
He repaid Rs. 7070
cost of goat = 17670-7070= Rs. 10,600